# algorithmic

### v1

@@ 2.3.1 Unsupervised tasks
Segmentation involves partitioning the data stream into contiguous subsequences or “blocks” $\{\mathcal{D}_{1:t_{1}},\mathcal{D}_{t_{1}+1:t_{2}},\ldots\}$, where the DGP for each block is governed by a sequence of unknown functions (Barry & Hartigan, 1992). The goal is to determine the points in time when a new block begins, known as changepoints. This is useful in many applications, such as finance, where detecting changes in market trends is critical. For further reference, see e.g, Aminikhanghahi & Cook (2017); Gupta et al. (2024). In this setting, non-stationarity is assumed to be abrupt and occurring at unknown points in time. We study an example in Section 4.3.1.
@@
Filtering aims to estimate an underlying latent state ${\bm{\theta}}_{t}$ which evolves over time and often represents a meaningful concept. The posterior estimate of ${\bm{\theta}}_{t}$ is computed by applying Bayesian inference to the corresponding state space model (SSM), which determines the choice of (M.1: likelihood), and how the state changes over time, through the choice of (M.3: prior). Examples include estimating the state of the atmosphere (Evensen, 1994), tracking the position of a moving object (Battin, 1982), or recovering a signal from a noisy system (Basseville et al., 1993). In this setting, non-stationarity is usually assumed to be continuous and occurring at possible time-varying rates.
@@ 2.3.2 Supervised tasks
Supervised tasks involve predicting a measurable outcome ${\bm{y}}_{t}$. Unlike unsupervised tasks, this allows the performance of the model to be assessed in an objective manner, since we can compare the prediction to the actual observation. Supervised tasks have been the main focus of the machine learning community. We present three common tasks in this category below.
@@
Prequential (or one-step-ahead) forecasting (Gama et al., 2008) seeks to predict the value ${\bm{y}}_{t+1}$ given $\mathcal{D}_{1:t}$ and ${\bm{x}}_{t+1}$. This is distinct from time-series forecasting, which typically does not consider exogenous variables ${\bm{x}}_{t}$, and thus can forecast (or “roll out”) many steps into the future. We study an example in Section 4.1.
@@
OCL is similar to prequential learning, in that the goal is to learn a supervised model for regression or classification online. However, the objective is to train a model that performs consistently across both past and future data, rather than just focusing on future forecasting (Cai et al., 2021). The changepoints (corresponding to different “tasks”) may or may not be known. This setting addresses the stability-plasticity dilemma, focusing on retaining previously learned knowledge while adapting to new tasks. We study an example of OCL for classification, when the task boundaries are not known, in Section 4.1.2.
@@
In contextual bandit problems, the agent is presented with features ${\bm{x}}_{t+1}$, and must choose an action (arm) that yields the highest expected reward (Li et al., 2010). We let ${\bm{y}}_{t+1}\in\mathbb{R}^{A}$ where $A>2$ is the number of possible actions; this is a vector where the $a$-th entry contains the reward one would have obtained had one chosen arm $a$. Let ${\bm{y}}_{t}^{(a)}$ be the observed reward at time $t$ after choosing arm $a$, i.e., the $a$-th entry of ${\bm{y}}_{t}$. A popular approach for choosing the optimal action (while tackling the exploration-exploitation tradeoff) at each step is Thomson sampling (TS) (Thompson, 1933), which in our setting works as follows: first, sample a parameter vector from the posterior, $\tilde{\bm{\theta}}_{t}$ from $\lambda({\bm{\theta}}_{t};\,\psi_{t},\mathcal{D}_{1:t})$; then, greedily choose the best arm (the one with the highest expected payoff), $a_{t+1}=\operatorname*{arg\,max}_{a}\hat{{\bm{y}}}_{t+1}^{(a)}$, where $\hat{\bm{y}}_{t+1}=h(\tilde{\bm{\theta}}_{t};{\bm{x}}_{t+1})$; and $\hat{{\bm{y}}}_{t+1}^{(a)}$ is the $a$-th entry of $\hat{{\bm{y}}}_{t+1}$; finally, receive a reward ${\bm{y}}_{t+1}^{(a_{t+1})}$. The goal is to select a sequence of arms $\{a_{1},\ldots,a_{T}\}$ that maximises the cumulative reward $\sum_{t=1}^{T}{\bm{y}}_{t}^{(a_{t})}$. When the mapping function $h$ is a neural network, this model is called a neural bandit. TS for neural bandits has been studied in many papers, see e.g., Duran-Martin et al. (2022) and references therein. Non-stationary bandits have been studied in Mellor & Shapiro (2013); Cartea et al. (2023a); Alami (2023); Liu et al. (2023). We study an example in Section 4.2.
@@ 2.4 Details of BONE
2.5 The measurement (likelihood) model (M.1)
Recall that $h({\bm{\theta}}_{t};\,{\bm{x}}_{t+1})$ is defined by
$h({\bm{\theta}}_{t};\,{\bm{x}}_{t+1})=\mathbb{E}[{\bm{y}}_{t+1}\,|\,{\bm{\theta}}_{t},{\bm{x}}_{t+1}].$ (7)
For example, for linear measurement models, the expected value is given by:
@@
RL (runlength): $\psi_{t}=r_{t}\in\{0,\ldots,t\}$ is a scalar representing a lookback window, defined as the number of steps since the last regime change. The value $r_{t}=0$ indicates the start of a new regime at time $t$, while $r_{t}\geq 1$ denotes the continuation of a regime with a lookback window of length $r_{t}$. This choice of auxiliary variable is common in the changepoint detection literature. See e.g., Adams & MacKay (2007); Knoblauch et al. (2018); Alami et al. (2020); Agudelo-España et al. (2020); Altamirano et al. (2023); Alami (2023).
RLCC (runlength and changepoint count): $\psi_{t}=(r_{t},c_{t})\in\{0,\ldots,t\}\times\{0,\ldots,t\}$ is a vector that represents both the runlength and the total number of changepoints, as proposed in Wilson et al. (2010). When $r_{t}=t$, this implies $c_{t}=0$, meaning no changepoints have occurred. Conversely, $r_{t}=0$ indicates the start of a new regime and implies $c_{t}\in\{1,\ldots,t\}$, accounting for at least one changepoint. For a given $r_{t}\geq 0$, the changepoint count $c_{t}$ belongs to the range $\{1,\ldots,t-r_{t}\}$. As with RL, this auxiliary variable assumes consecutive time blocks, but additionally allows us to estimate the likelihood of entering a new regime by tracking the number of changepoints seen so far.
CPT (changepoint timestep): $\psi_{t}=\boldsymbol{\zeta}_{t}$, with $\boldsymbol{\zeta}_{t}=\{\zeta_{1,t},\ldots,\zeta_{\ell,t}\}$, is a set of size $\ell\in\{0,\ldots t\}$ containing the $\ell$ times at which there was a changepoint, with the convention that $0\leq\zeta_{1,t}<\zeta_{2,t}<\ldots<\zeta_{\ell,t}\leq t$. This choice of auxiliary variable was introduced in Fearnhead & Liu (2007) and has been studied in Fearnhead & Liu (2011); Fearnhead & Rigaill (2019). Under mild assumptions, it can be shown that CPT is equivalent to RL, see e.g., Knoblauch & Damoulas (2018).
CPL (changepoint location): $\psi_{t}=s_{1:t}\in\{0,1\}^{t}$ is a binary vector. In one interpretation, $s_{i}=1$ indicates the occurrence of a changepoint at time $i$, as in Li et al. (2021), while in another, it means that $\mathcal{D}_{t}$ belongs to the current regime, as in Nassar et al. (2022).
CPV (changepoint probability vector): $\psi_{t}=v_{1:t}\in(0,1)^{t}$ is a $t$-dimensional random vector representing the probability of each element in the history belonging to the current regime. This generalises CPL and was introduced in Nassar et al. (2022) for online continual learning, allowing for a more fine-grained representation of changepoints over time.
@@
ME (mixture of experts): $\psi_{t}=\alpha_{t}\in\{1,\ldots,K\}$ represents one of $K$ experts. Each expert corresponds to either a choice of model or one of $K$ possible hyperparameters. This approach has been applied to filtering (Chaer et al., 1997) and prequential forecasting (Liu, 2023; Abélès et al., 2024).
C: $\psi_{t}=c$ represents a constant auxiliary variable, where $c$ is just a placeholder or dummy value. This is equivalent to not having an auxiliary variable, or alternatively, to having a single expert that encodes all available information; this choice recovers the stationary model in (1).
@@
values $\{0\}$ $2^{\{0,1,\ldots,t\}}$ $[0,1]$ $\{0,1\}^{t}$ $(0,1)^{t}$$\{1,\ldots,K\}$$\{0,1,\ldots,t\}$$\{\{0,t\},\ldots,\{t,0\}\}$
@@
This component defines the prior over model parameters conditioned on the choice of (M.2: auxvar) $\psi_{t}$ and the dataset $\mathcal{D}_{1:t-1}$. In some cases, explicit access to past data is not needed.
@@
$\tau({\bm{\theta}}_{t}\,|\,\psi_{t},\,\mathcal{D}_{1:t-1})={\cal N}\big{(}{\bm{\theta}}_{t}\,|\,g_{t}(\psi_{t},\mathcal{D}_{1:t-1}),G_{t}(\psi_{t},\mathcal{D}_{1:t-1})\big{)},$ (9)
with $g_{t}:\boldsymbol{\Psi}_{t}\times^{(m+d)(t-1)}\to^{m}$ a function that returns the mean vector of model parameters, and $G_{t}:\boldsymbol{\Psi}_{t}\times^{(m+d)(t-1)}\to^{m\times m}$ a function that returns a $m$-dimensional covariance matrix.
@@
C-LSSM (constant linear with affine state-space model). We assume the parameter dynamics can be modeled by a linear-Gaussian state space model (LSSM), i.e., $\mathbb{E}[{\bm{\theta}}_{t}\,|\,{\bm{\theta}}_{t-1}]=\mathbf{F}_{t}\,{\bm{\theta}}_{t-1}+{\bm{b}}_{t}$ and ${\rm Var}({\bm{\theta}}_{t})=\mathbf{Q}_{t}$, for given $(m\times m)$ dynamics matrix $\mathbf{F}_{t}$, $(m\times 1)$ bias vector ${\bm{b}}_{t}$, and $(m\times m)$ positive semi-definite matrix $\mathbf{Q}_{t}$. We also assume $\psi_{t}=c$ is a fixed (dummy) constant, which is equivalent to not having an auxiliary variable. The characterisation of the conditional prior takes the form
@@
$G_{t}(c,\mathcal{D}_{1:t-1})$ $=\mathbf{F}_{t}\,\boldsymbol{\Sigma}_{t-1}\mathbf{F}_{t}^{\intercal}+\mathbf{Q}_{t}\,.$
@@
C-OU (constant with Ornstein-Uhlenbeck process). This is a special case of the C-LSSM model where $\mathbf{F}_{t}=\gamma\mathbf{I}$, ${\bm{b}}_{t}=(1-\gamma)\boldsymbol{\mu}_{0}$, and $\mathbf{Q}_{t}=(1-\gamma^{2})\boldsymbol{\Sigma}_{0}$, where $\gamma$ is the fixed rate. The conditional prior mean and covariance are a convex combination of the form
@@
CPP-OU (changepoint probability with Ornstein-Uhlenbeck process). Here $\psi_{t}=\upsilon_{t}\in(0,1)$ is the changepoint probability that we use as the rate of an Ornstein-Uhlenbeck (OU) process. This combination is used in Titsias et al. (2024); Galashov et al. (2024). The characterisation of the conditional prior takes the form
@@
This is similar to C-OU with $\gamma=1$, however, here we inject new noise at each step.
@@
RL[1]-OUPR* (greedy runlength with OU and prior reset): This is a new combination we consider in this paper, which is designed to accommodate both gradual changes and sudden changes. More precisely, we assume $\psi_{t}=r_{t}$, and we choose the conditional prior as either a hard reset to the prior (if the posterior weight $\nu(r_{t})=p(r_{t}\,|\,\mathcal{D}_{1:t})$, for $r_{t}\geq 1$, is below a threshold $\varepsilon$), or a convex combination of the prior and the previous belief state (corresponding to a choice of OU). We characterise the conditional prior as
$g_{t}(r_{t},\mathcal{D}_{1:t-1})=\begin{cases}\boldsymbol{\mu}_{0}\,(1-\nu(r_{t}))+\boldsymbol{\mu}_{(r_{t})}\,\nu(r_{t})&\nu(r_{t})>\varepsilon,\\ \boldsymbol{\mu}_{0}&\nu(r_{t})\leq\varepsilon,\end{cases}$ (17)
$G_{t}(r_{t},\mathcal{D}_{1:t-1})=\begin{cases}\boldsymbol{\Sigma}_{0}\,(1-\nu(r_{t})^{2})+\boldsymbol{\Sigma}_{(r_{t})}\,\nu(r_{t})^{2}&\nu(r_{t})>\varepsilon,\\ \boldsymbol{\Sigma}_{0}&\nu(r_{t})\leq\varepsilon.\end{cases}$ (18)
Here $\nu(r_{t})\in[0,1]$ is a function of $r_{t}$ and corresponds to the choice of (A.2: weighting); see Section 2.9. The value of the threshold parameter $\varepsilon$ controls whether an abrupt change or a gradual change should take place. In the limit when $\varepsilon=1$, this new combination does not learn and is always at the prior beliefs. Conversely, when $\varepsilon=0$, we obtain an OU-type update weighted by $\nu$. When $\varepsilon=0.5$, we revert back to prior beliefs when the most likely hypothesis is that a changepoint has just occurred. Finally, we remark that the above combination allows us to make use of non-Markovian choices for (M.1: likelihood), as we see in Section 4.3.1. This is, to the best of our knowledge, a new combination that has not been proposed in the previous literature; for further details see Appendix A.3.
@@
This section presents algorithms for estimating the density $\lambda({\bm{\theta}}_{t};\,\psi_{t},\mathcal{D}_{1:t})$; we focus on methods that yield Gaussian posterior densities. Specifically, we are interested in practical approaches for approximating the conditional Bayesian posterior, as given in (4).
There is a vast body of literature on methods for estimating the posterior over model parameters. Here, we focus on three common approaches for computing the Gaussian posterior: conjugate updates (Cj), linear-Gaussian approximations (LG), and variational Bayes (VB).

### v3

@@ 2.1.1 Unsupervised tasks
Segmentation involves partitioning the data stream into contiguous subsequences or “blocks” $\{\mathcal{D}_{1:t_{1}},\mathcal{D}_{t_{1}+1:t_{2}},\ldots\}$, where the DGP for each block is governed by a sequence of unknown functions (Barry & Hartigan, 1992). The goal is to determine the points in time when a new block begins, known as changepoints. This is useful in many applications, such as finance, where detecting changes in market trends is critical. In this setting, non-stationarity is assumed to be abrupt and occurring at unknown points in time. We study an example in Section 4.3.1. For a survey of segmentation methods, see e.g, Aminikhanghahi & Cook (2017); Gupta et al. (2024).
@@
Filtering estimates an underlying latent state ${\bm{\theta}}_{t}$ that evolves over time (often representing a meaningful concept). The posterior estimate of ${\bm{\theta}}_{t}$ is computed by applying Bayesian inference to the corresponding state space model (SSM), which determines the choice of (M.1: model), and how the state changes over time, through the choice of (M.3: prior). Examples include estimating the state of the atmosphere (Evensen, 1994), tracking the position of a moving object (Battin, 1982), or recovering a signal from a noisy system (Basseville et al., 1993). In this setting, non-stationarity is usually assumed to be continuous and occurring at possible time-varying rates. For a survey of filtering methods, see e.g., Chen et al. (2003).
@@ 2.1.2 Supervised tasks
Supervised tasks involve predicting a measurable outcome ${\bm{y}}_{t}$. Unlike unsupervised tasks, this allows the performance of the model to be assessed objectively, since we can compare the prediction to the actual observation. We present three common tasks in this category below.
@@
Prequential (or one-step-ahead) forecasting (Gama et al., 2008) seeks to predict the value ${\bm{y}}_{t+1}$ given $\mathcal{D}_{1:t}$ and ${\bm{x}}_{t+1}$. This is distinct from time-series forecasting, which typically does not consider exogenous variables ${\bm{x}}_{t}$, and thus can forecast (or “roll out”) many steps into the future. We study an example in Section 4.1. For a survey on prequential forecasting under non-stationarity, see e.g., Lu et al. (2018).
@@
OCL is a broad term used for learning regression or classification models online, typically with neural networks. These methods usually assume that the underlying data generating mechanism could shift. The objective of OCL methods is to train a model that performs consistently across both past and future data, rather than just focusing on future forecasting (Cai et al., 2021). The changepoints (corresponding to different “tasks”) may or may not be known. This setting addresses the stability-plasticity dilemma, focusing on retaining previously learned knowledge while adapting to new tasks. We study an example of OCL for classification, when the task boundaries are not known, in Section 4.1.2. For a survey on recent methods for OCL, see e.g., Gunasekara et al. (2023).
@@
In contextual bandit problems, the agent is presented with features ${\bm{x}}_{t+1}$, and must choose an action (arm) that yields the highest expected reward (Li et al., 2010). We let ${\bm{y}}_{t+1}\in\mathbb{R}^{A}$ where $A>2$ is the number of possible actions; this is a vector where the $a$-th entry contains the reward one would have obtained had one chosen arm $a$. Let ${\bm{y}}_{t}^{(a)}$ be the observed reward at time $t$ after choosing arm $a$, i.e., the $a$-th entry of ${\bm{y}}_{t}$. A popular approach for choosing the optimal action (while tackling the exploration-exploitation tradeoff) at each step is Thomson sampling (TS) (Thompson, 1933), which in our setting works as follows: first, sample a parameter vector from the posterior, $\tilde{\bm{\theta}}_{t}$ from $q({\bm{\theta}}_{t};\,\psi_{t},\mathcal{D}_{1:t})$; then, greedily choose the best arm (the one with the highest expected payoff), $a_{t+1}=\operatorname*{arg\,max}_{a}\hat{{\bm{y}}}_{t+1}^{(a)}$, where $\hat{\bm{y}}_{t+1}=h(\tilde{\bm{\theta}}_{t};{\bm{x}}_{t+1})$; and $\hat{{\bm{y}}}_{t+1}^{(a)}$ is the $a$-th entry of $\hat{{\bm{y}}}_{t+1}$; finally, receive a reward ${\bm{y}}_{t+1}^{(a_{t+1})}$. The goal is to select a sequence of arms $\{a_{1},\ldots,a_{T}\}$ that maximises the cumulative reward $\sum_{t=1}^{T}{\bm{y}}_{t}^{(a_{t})}$. When the mapping function $h$ is a neural network, this model is called a neural bandit. TS for neural bandits has been studied in many papers, see e.g., Duran-Martin et al. (2022) and references therein. Non-stationary bandits have been studied in Mellor & Shapiro (2013); Cartea et al. (2023a); Alami (2023); Liu et al. (2023). We study an example in Section 4.2.
@@ 2.2 Details of BONE
2.3 The measurement model (M.1)
Recall that $h({\bm{\theta}},{\bm{x}})$ is a parametric model that encodes the conditional mean for ${\bm{y}}$, given ${\bm{\theta}}$ and ${\bm{x}}$. For linear measurement models, $h({\bm{\theta}},{\bm{x}})$ is given by:
@@
RL (runlength): $\psi_{t}=r_{t}\in\{0,\ldots,t\}$ is a scalar representing a lookback window, defined as the number of steps since the last regime change. The value $r_{t}=0$ indicates the start of a new regime at time $t$, while $r_{t}\geq 1$ denotes the continuation of a regime with a lookback window of length $r_{t}$. This choice of auxiliary variable is common in the changepoint detection literature. See e.g., Adams & MacKay (2007); Knoblauch et al. (2018); Alami et al. (2020); Agudelo-España et al. (2020); Altamirano et al. (2023); Alami (2023). This auxiliary variable is useful for non-stationary data with non-repeating temporal segments, provided we know the intensity with which new segments appear.
RLCC (runlength and changepoint count): $\psi_{t}=(r_{t},c_{t})\in\{0,\ldots,t\}\times\{0,\ldots,t\}$ is a vector that represents both the runlength and the total number of changepoints, as proposed in Wilson et al. (2010). When $r_{t}=t$, this implies $c_{t}=0$, meaning no changepoints have occurred. Conversely, $r_{t}=0$ indicates the start of a new regime and implies $c_{t}\in\{1,\ldots,t\}$, accounting for at least one changepoint. For a given $r_{t}\geq 0$, the changepoint count $c_{t}$ belongs to the range $\{1,\ldots,t-r_{t}\}$. As with RL, this auxiliary variable assumes consecutive time blocks, but additionally allows us to estimate the likelihood of entering a new regime by tracking the number of changepoints seen so far. This auxiliary variable is useful for non-stationary data with non-repeating temporal segments when the intensity with which new segments appear is unknown.
CPT (changepoint timestep): $\psi_{t}=\boldsymbol{\zeta}_{t}$, with $\boldsymbol{\zeta}_{t}=\{\zeta_{1,t},\ldots,\zeta_{\ell,t}\}$, is a set of size $\ell\in\{0,\ldots t\}$ containing the $\ell$ times at which there was a changepoint, with the convention that $0\leq\zeta_{1,t}<\zeta_{2,t}<\ldots<\zeta_{\ell,t}\leq t$. This choice of auxiliary variable was introduced in Fearnhead & Liu (2007) and has been studied in Fearnhead & Liu (2011); Fearnhead & Rigaill (2019). Under mild assumptions, it can be shown that CPT is equivalent to RL, see e.g., Knoblauch & Damoulas (2018). This auxiliary variable is useful for non-stationary data with non-repeating temporal segments when the probability of a new segment appearing is unknown and knowledge of the changepoint location is required.
CPL (changepoint location): $\psi_{t}=s_{1:t}\in\{0,1\}^{t}$ is a binary vector. In one interpretation, $s_{i}=1$ indicates the occurrence of a changepoint at time $i$, as in Li et al. (2021), while in another, it means that $\mathcal{D}_{t}$ belongs to the current regime, as in Nassar et al. (2022). This auxiliary variable is useful for non-stationary data with repeating temporal segments. It is useful when the segments are formed of non-consecutive datapoints.
CPV (changepoint probability vector): $\psi_{t}=v_{1:t}\in(0,1)^{t}$ is a $t$-dimensional random vector representing the probability of each element in the history belonging to the current regime. This generalises CPL and was introduced in Nassar et al. (2022) for online continual learning, allowing for a more fine-grained representation of changepoints over time. This auxiliary variable is useful for non-stationary data with repeating temporal segments. Unlike CPL, it takes a vector of weights in $(0,1)$ which allows for higher flexibilty when compared to CPL.
@@
ME (mixture of experts): $\psi_{t}=\alpha_{t}\in\{1,\ldots,K\}$ represents one of $K$ experts. Each expert corresponds to either a choice of model or one of $K$ possible hyperparameters. This approach has been applied to filtering (Chaer et al., 1997) and prequential forecasting (Liu, 2023; Abélès et al., 2024). This auxiliary variable facilitates the weighting of predictions made by models when one has a fixed number of competing models.
C: $\psi_{t}=c$ represents a constant auxiliary variable, where $c$ is just a placeholder or dummy value. This is equivalent to not having an auxiliary variable, or alternatively, to having a single expert that encodes all available information.
@@
values $\{c\}$ $2^{\{0,1,\ldots,t\}}$ $[0,1]$ $\{0,1\}^{t}$ $(0,1)^{t}$$\{1,\ldots,K\}$$\{0,1,\ldots,t\}$$\{\{0,t\},\ldots,\{t,0\}\}$
@@
This component defines the prior predictive distribution over model parameters conditioned on the choice of (M.2: auxvar) $\psi_{t}$ and the dataset $\mathcal{D}_{1:t-1}$. In some cases, explicit access to past data is not needed.
@@
$\pi({\bm{\theta}}_{t};\,\psi_{t},\,\mathcal{D}_{1:t-1})={\cal N}\big{(}{\bm{\theta}}_{t}\,|\,g_{t}(\psi_{t},\mathcal{D}_{1:t-1}),G_{t}(\psi_{t},\mathcal{D}_{1:t-1})\big{)},$ (11)
with $g_{t}:\boldsymbol{\Psi}_{t}\times^{(m+d)(t-1)}\to^{m}$ a function that returns the mean vector of model parameters, $\mathbb{E}[{\bm{\theta}}_{t}\,|\,\psi_{t},\mathcal{D}_{1:t-1}]$, and $G_{t}:\boldsymbol{\Psi}_{t}\times^{(m+d)(t-1)}\to^{m\times m}$ a function that returns a $m$-dimensional covariance matrix, ${\rm Cov}[{\bm{\theta}}_{t}\,|\,\psi_{t},\mathcal{D}_{1:t-1}]$. In what follows, we let $(\boldsymbol{\mu}_{0},\boldsymbol{\Sigma}_{0})$ be the pre-defined initial prior mean and covariance. Furthermore, we denote $(\boldsymbol{\mu}_{t-1},\boldsymbol{\Sigma}_{t-1})$ be the posterior mean and covariance found at time $t-1$, which is used as a prior at time $t$.
@@
C-LSSM (constant linear with affine state-space model). We assume the parameter dynamics can be modeled by a linear-Gaussian state space model (LSSM), i.e., $\mathbb{E}[{\bm{\theta}}_{t}\,|\,{\bm{\theta}}_{t-1}]=\mathbf{F}_{t}\,{\bm{\theta}}_{t-1}+{\bm{b}}_{t}$ and ${\rm Cov}[{\bm{\theta}}_{t}\,|\,{\bm{\theta}}_{t-1}]=\mathbf{Q}_{t}$, for given $(m\times m)$ dynamics matrix $\mathbf{F}_{t}$, $(m\times 1)$ bias vector ${\bm{b}}_{t}$, and $(m\times m)$ positive semi-definite matrix $\mathbf{Q}_{t}$. We also assume $\psi_{t}=c$ is a fixed (dummy) constant, which is equivalent to not having an auxiliary variable. The characterisation of the conditional prior takes the form
@@
$G_{t}(c,\mathcal{D}_{1:t-1})$ $=\mathbf{F}_{t}\,\boldsymbol{\Sigma}_{t-1}\mathbf{F}_{t}^{\intercal}+\mathbf{Q}_{t}\,,$
@@
C-OU (constant with Ornstein-Uhlenbeck process). This is a special case of the C-LSSM model where $\mathbf{F}_{t}=\gamma\mathbf{I}$, ${\bm{b}}_{t}=(1-\gamma)\boldsymbol{\mu}_{0}$, $\mathbf{Q}_{t}=(1-\gamma^{2})\boldsymbol{\Sigma}_{0}$, $\boldsymbol{\Sigma}_{0}=\sigma_{0}^{2}\mathbf{I}$, $\gamma\in[0,1]$ is the fixed rate, and $\sigma_{0}\geq 0$. The conditional prior mean and covariance are a convex combination of the form
@@
CPP-OU (changepoint probability with Ornstein-Uhlenbeck process). Here $\psi_{t}=\upsilon_{t}\in[0,1]$ is the changepoint probability that we use as the rate of an Ornstein-Uhlenbeck (OU) process, as proposed in Titsias et al. (2024); Galashov et al. (2024). The characterisation of the conditional prior takes the form
@@
An example on how to compute $\upsilon_{t}$ using an empirical Bayes procedure is given in (42).
@@
This is similar to C-OU with $\gamma=1$, however, here we inject new noise at each step. Another variant of this scheme, known as shrink-and-perturb (Ash & Adams, 2020), takes $g(c,\mathcal{D}_{1:t-1})=q\,\boldsymbol{\mu}_{t-1}$ and $G(c,\mathcal{D}_{1:t-1})=\boldsymbol{\Sigma}_{t-1}+\mathbf{Q}_{t}$, where $10<q<$ 1 is the shrinkage parameters, and $\mathbf{Q}_{t}=\sigma_{0}^{2}\,\mathbf{I}$.
@@
RL[1]-OUPR* (greedy runlength with OU and prior reset): This is a new combination we consider in this paper, which is designed to accommodate both gradual changes and sudden changes. More precisely, we assume $\psi_{t}=r_{t}$, and we choose the conditional prior as either a hard reset to the prior, if $\nu_{t}(r_{t})>\varepsilon$, or a convex combination of the prior and the previous belief state (using an OU process), if $\nu_{t}(r_{t})\leq\varepsilon$. That is, we define the conditional prior as
$g_{t}(r_{t},\mathcal{D}_{1:t-1})=\begin{cases}\boldsymbol{\mu}_{0}\,(1-\nu_{t}(r_{t}))+\boldsymbol{\mu}_{(r_{t})}\,\nu_{t}(r_{t})&\nu_{t}(r_{t})>\varepsilon,\\ \boldsymbol{\mu}_{0}&\nu_{t}(r_{t})\leq\varepsilon,\end{cases}$ (19)
$G_{t}(r_{t},\mathcal{D}_{1:t-1})=\begin{cases}\boldsymbol{\Sigma}_{0}\,(1-\nu_{t}(r_{t})^{2})+\boldsymbol{\Sigma}_{(r_{t})}\,\nu_{t}(r_{t})^{2}&\nu_{t}(r_{t})>\varepsilon,\\ \boldsymbol{\Sigma}_{0}&\nu_{t}(r_{t})\leq\varepsilon.\end{cases}$ (20)
Here $\nu_{t}(r_{t})=p(r_{t}\,|\,\mathcal{D}_{1:t})$, with $r_{t}=r_{t-1}+1$, is the probability we are continuing a segment, and $\nu_{t}(r_{t})$ with $r_{t}=0$ is the probability of a changepoint. For details on how to compute $\nu_{t}(r_{t})$, see (37). The value of the threshold parameter $\varepsilon$ controls whether an abrupt change or a gradual change should take place. In the limit when $\varepsilon=1$, this new combination does not learn, since it is always doing a hard reset to the initial beliefs at time $t=0$. Conversely, when $\varepsilon=0$, we obtain an OU-type update weighted by $\nu_{t}$. When $\varepsilon=0.5$, we revert back to prior beliefs when the most likely hypothesis is that a changepoint has just occurred. Finally, we remark that the above combination allows us to make use of non-Markovian choices for (M.1: model), as we see in Section 4.3.1. This is, to the best of our knowledge, a new combination that has not been proposed in the previous literature; for further details see Appendix A.3.
@@
This section presents algorithms for estimating the density $q({\bm{\theta}}_{t};\,\psi_{t},\mathcal{D}_{1:t})$; we focus on methods that yield Gaussian posterior densities. Specifically, we are interested in practical approaches for approximating the conditional Bayesian posterior, as given in (7).
There is a vast body of literature on methods for estimating the posterior over model parameters. Here, we focus on three common approaches for computing the Gaussian posterior: conjugate updates (Cj), linear-Gaussian approximations (LG), and variational Bayes (VB). For an overview of choices of (A.1: posterior) and a comparison in terms of their computational complexity, see Table 3 in Jones et al. (2024).